OpenStudy (anonymous):

the ratio of men to women in a certain factory is 3 to 4. there are 210 men. how many workers are there.?

OpenStudy (anonymous):

can any one work out this problem?

OpenStudy (anonymous):

m/w = 3/4 m = 210 w=?

OpenStudy (anonymous):

the total number is X, WHOLE FRACTION IN A PIE DIAGRAM is 3 + 4 =7 , so if 3/7 are men, = 210, it can be written as 3/7 of x = 210, what X?

OpenStudy (anonymous):

i need the answer for W?

OpenStudy (anonymous):

well, just solve for x, as I have said, then, that will be the total number of people in the factory. to get the number of women, W, = M + W =X 210 + W =X if you find the X value, then you will find w value